OpenStudy (anonymous):

The vertex of the function’s graph is given. Find c. y = c - x2 - 2x; (-1, 3) a. 5 c. -3 b. -2 d. 2

OpenStudy (anonymous):

d. Just substitute x=-1 and y=3 into the eqn

OpenStudy (anonymous):

3 = c - (-1)^2 - 2(-1) 3 = c - 1+2 3= c +1 c=2